# silent-chen.github.io/DGE

### v1

@@ Abstract
We consider the problem of editing 3D objects and scenes based on open-ended language instructions. The established paradigm to solve this problem is to use a 2D image generator or editor to guide the 3D editing process. However, this is often slow as it requires do update a computationally expensive 3D representations such as a neural radiance field, and to do so by using contradictory guidance from a 2D model which is inherently not multi-view consistent. We thus introduce the Direct Gaussian Editor (DGE), a method that addresses these issues in two ways. First, we modify a given high-quality image editor like InstructPix2Pix to be multi-view consistent. We do so by utilizing a training-free approach which integrates cues from the underlying 3D geometry of the scene. Second, given a multi-view consistent edited sequence of images of the object, we directly and efficiently optimize the 3D object representation, which is based on 3D Gaussian Splatting. Because it does not require to apply edits incrementally and iteratively, DGE is significantly more efficient than existing approaches, and comes with other perks such as allowing selective editing of parts of the scene.
@@ 1 Introduction
AI-based content creation and editing have advanced substantially due to recent breakthroughs in 2D and 3D generation [82, 33, 78]. The benefits are particularly significant for 3D content, processing which normally requries highly-specialized talent and software. AI-powered tools are unlocking new creative possibilities for artists and non-professional users alike.
@@
In this work, we consider the problem of editing 3D models based on textual instructions. Part of the recent progress in this area is due to the introduction of new radiance field representations [69, 43] for objects and scenes, which are both high-quality and flexible. Because they are differentiable, radiance fields combine well with recent advances in image generation. A notable example is InstructPix2Pix (IP2P) [4], a denoising diffusion model [89, 83] that can edit the content of an image based on textual instructions [4]. This and similar models can be used to update radiance fields, thus resulting in open-ended editing of 3D contents. So far, this has been done primarily in two ways. The first was pioneered by InstructNeRF2NeRF [28], which utilizes IP2P to alternate between editing one rendered view of the 3D model and updating the latter. The second approach [123, 40, 85] is to update the model in a distillation-like fashion [78]. Either way, edits are slowly incorporated by alternating between updating the 3D model and generating and processing different views of it, so that a single edit often requires several minutes, and sometimes even hours, to complete.
@@
In this paper, we introduce a new approach to 3D editing, the Direct Gaussian Editor (DGE), which addresses this and other shortcomings of these earlier attempts. We design our method with three goals in mind: (i) high fidelity, (ii) high efficiency, and (iii) selective editing (i.e., the ability to only edit part of an object or a scene).
@@
To achieve these goals, we propose to change both the representation of the 3D model and the way it is updated. As a representation, we use Gaussian Splatting (GS) [43]. This is an alternative representation of a radiance field with the main obvious advantage of being orders of magnitude faster than NeRF and NeRF-adjacent models [11, 91, 9], for both rendering and gradient computation. There is another, less obvious but also important advantage of GS: because it is an explicit 3D representation made of local 3D primitives, the Gaussians, it supports easily and efficiently local edits, as long as one knows which Gaussian(s) to modify. In practice, we can easily identify the Gaussians by fusing the output of a 2D instance segmenter [45] from several rendered views of the scene.
While GS can contribute to the speed of the 3D editor, it does not remove the bottleneck of iterating several times between rendering, reconstruction and evaluation of the underlying image-based diffusion model, meaning that edits would still require a considerable amount of time to complete. Hence, our second contribution is to develop a faster alternative to this slow iterative process.
@@
The reason why distillation-like methods are used in prior works is the lack of a set of view-consistent edits to update the 3D model coherently. Because 2D editors like IP2P are monocular and provide a distribution over all possible edits, the probability of drawing independently two or more consistent edits is nearly zero. Distillation is a way to incrementally reach a “multi-view” consensus w.r.t the application of the monocular editor to the different views.
@@
Here, we suggest to use instead a method that can sample multi-view consistent edits such that the 3D model can be updated by directly fitting it to the edited views. Our approach to multi-view consistent editing is inspired by recent progress in video generation and editing and, in particular, methods that extend image generators to video without additional training [23, 106, 44, 8, 80]. The key insight is that multiple views of the 3D model can be interpreted as a video of a static scene generated by a moving camera. This then means that achieving multi-view consistency is analogous to temporal consistency in video editing. To achieve multi-view consistency we adopt the spatio-temporal attention mechanism used in the video editing literature and extend it with additional epipolar constraints, which we can enforce due to the 3D nature of our problem.
@@
Through qualitative and quantitative comparisons to prior works, we demonstrate two key advantages of DGE, even when compared to very recent works like the GaussianEditor [15], which also use Gaussian Splatting. The first advantage is that, as a direct editing approach, DGE is more efficient and results in a noticeable speed boost compared to prior approaches (approximately 4mins for a single edit). Second, ensuring multi-view consistent editing in the image space significantly simplifies the process of integrating edits from various views into the 3D model. This is then reflected in both the number of updates that are required for convergence as well as the higher fidelity of the final result.
@@ 2 Related Work
Most methods for 3D image editing are based on 2D image editing due to the lack of large scale training data to learn 3D editors directly, so we discuss those first. GLIDE [74] controls 2D image generation and editing using CLIP features [81]. Methods like [21, 49] and DreamBooth [84] consider the problem of personalizing an image; GLIGEN [57] and others [10, 13, 20] perform layout control and [118] considers other forms of control. DragDiffusion [86] and Drag-a-Video [93] edit images and videos by dragging. Others [65, 72, 2, 96, 76, 42, 31] cast image editing as image-to-image translation. InstructPix2Pix (IP2P) [5] fine-tunes Stable Diffusion [82] large synthetic dataset of example language-driven edits. The works of [117, 120] further improve InstructPix2Pix via manual labelling and the work of [70] aims at localizing the edits better.
Ad-hoc 3D Editing.
@@
Several authors have explored various types of inputs and controls to edit 3D objects. EditNeRF [60] updates shape and color in a radiance field based on user-provided scribbles, 3Designer [54] and SINE [1] based on a single edited view, SKED [67] based on 2D sketches, Editable-NeRF [121] based on keypoints and CoNeRF [41] based on attributes. The work of [61] also considers sketch-based editing, but for category-level 3D generators. NeRF-Editing [113], NeuMesh [110] and NeRFShop [39] modify a radiance field based on meshes and the work of [109] based on cages. NeuralEditor [12] does so based on point clouds. N3F [95], DFF [46], SPIn-NeRF [71], NeRF-in [59] and the work of [105] consider segmenting and removing objects from a radiance field. Control-NeRF [50] allows both removing and moving objects. Component-NeRF [58] and the works of [115, 111] consider compositional editing. Palette-NeRF [48], RecolorNeRF [25], ICE-NeRF [51] address recoloring and ARF [116], DeSRF [108], StylizedNeRF [37], SNeRF [73] and the work of [18, 36] style transfer. NeRFEditor [92] integrates GAN-based stylization in NeRF. Seal-3D [104] consider pixel-level editing of 3D scenes learning interactive tools like brushes, deformation and recoloring and SealD-NeRF [38] extend it to dynamic NeRF models. SceNeRFlow [94] allows for dynamic edits as well using DensePose for correspondence estimation [27].
Open-ended 3D Editing.
@@
Like us, several authors have considered open-ended and language-based 3D editing. Some authors start from CLIP or other vision-language model [62, 53, 55] to drive editing/stylisation of a radiance field as a whole, like CLIP-NeRF [101], NeRF-Art [102], Text2Mesh [66] and Tango [52], or in part [87, 47, 103, 26]. Some like AvatarCLIP [35] consider animations as well. Most recent works start instead from diffusion-based image generators like StableDiffusion [82] or editors like Instruct-Pix2Pix [4]. Most of these methods are iterative. Some, like Instruct-NeRF2NeRF [29], InstructP2P [107], ProteusNeRF [100], Edit-DiffNeRF [112], Vica-NeRF [19] edit views of the object and update the 3D model accordingly. Others, like Instruct3Dto3D [40], DreamEditor [123], Vox-E [85], ED-NeRF [75], FocalDreamer [56], Progressive3D [17] and [114, 122], do the same, but using score distillation sampling. TextDeformer [22] focuses on manipulating only the shape of objects. Blending-NeRF [88] blends an edited radiance field with the original one. SHAP-EDITOR [14] learns to quickly edit 3D shapes encoded in a latent space, but require retraining for each new set of instructions.
@@
Our method allows direct editing, or substantially reduces the number of iterations required. Furthermore, methods like Vox-E [85] and FocalDreamer [56] consider local 3D editing, but struggle to handle both local and global ones; in contrast, our DGE allows precise editing of small or large regions of the 3D scene.
@@ 3 Preliminaries
Gaussian Radiance Fields.
A radiance field is a pair of functions $\sigma:\mathbb{R}^{3}\rightarrow\mathbb{R}_{+}$ and $c:\mathbb{R}^{3}\times\mathbb{S}^{2}\rightarrow\mathbb{R}^{3}$ mapping 3D points $\boldsymbol{x}\in\mathbb{R}^{3}$ to opacities $\sigma(\boldsymbol{x})$ and directional colors $c(\boldsymbol{x},\boldsymbol{\nu})$, where $\boldsymbol{\nu}\in\mathbb{S}^{2}$ is a unit vector expressing the viewing direction. to obtain an image $I$ of the radiance field, one uses the emission absorption equation:
@@
Prior works have explored several representations for these functions, including MLPs [69], voxel grids [91] and low-rank factorizations of the latter [11, 9]. Gaussian Splatting (GS) [43] proposes yet another but very efficient representation. GS is given by a mixture of Gaussians $\mathcal{G}=\{(\sigma_{i},\boldsymbol{\mu}_{i},\Sigma_{i},c_{i})\}_{i=1}^{G}$, where $\sigma_{i}\geq 0$ is the opacity, $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ is the mean, $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix, and $c_{i}:\mathbb{S}\rightarrow\mathbb{R}^{3}$ the directional color of each Gaussian. The Gaussian functions are given by:
@@
Diffusion-based Generators and Editors.
A diffusion-based image generator is a model that allows one to sample image $I\in\mathbb{R}^{3\times H\times W}$ from a conditional distribution $p(I\mid y)$. Here the condition $y$ is often a textual prompt, but it can also be an image or a combination of both. Let $I_{k}=\sqrt{1-\sigma_{k}^{2}}I+\sigma_{k}\epsilon$ be a noised version of the image $I$ one wishes to sample, where $\sigma$ is a sequence monotonically increasing from 0 to 1 and $\epsilon\sim\mathcal{N}$ is normal i.i.d. noise. The diffusion model consists of a denoising neural network $\Phi$ that, given the noisy image $I_{k}$, the noise level $k$ and the conditioning $y$, estimates the noise $\hat{\epsilon}=\Phi(I,y,k).$ The image $I$ is sampled by starting from pure noise ($\sigma_{T}=1$) followed by iterative denoising of the signal until $I=I_{0}$ is obtained [34, 89, 98].
Several authors have suggested to use the denoising network of an off-the-shelf image generator such as Stable Diffusion (SD) [83, 77] to modify an existing image according to a prompt $y$ instead of generating it from scratch, thus editing it. However, this process is somewhat suboptimal; instead, we make use of editors specifically trained to apply such modifications. An example is IP2P [6], which also uses diffusion but directly implements conditional distribution $p(I^{\prime}\mid I,y)$ where $I^{\prime}$ is the edited image, $I$ is the original image, and $y$ is a prompt describing how $I^{\prime}$ should be obtained from $I$. IP2P is in itself derived from the SD model by finetuning it to training triplets $(I^{\prime},I,y)$ which are automatically generated using Prompt-to-Prompt [32], a slow but training-free editing method.
@@ 4 Method
In this section, we describe our method for text-guided 3D scene editing — the Direct Gaussian Editor (DGE). Our main objective is 3D editing that is (i) faithful to the text instructions (high fidelity), (ii) fast and efficient, and (iii) partial to specific scene elements when desired. Next, we introduce our approach and discuss how these characteristics are met.
@@ 4.1 Direct Gaussian Editor
DGE is motivated by the following key observation: if one can obtain multi-view consistent edits in image space, then the 3D model can be updated directly and efficiently by fitting it to the edited images. This approach is alternative to slow iterative techniques like SDS [79] and iterative dataset updates (IDU) [28] that work around inconsistent image-level edits. Thus, the main challenge to allow direct 3D editing is to improve image-based editors to produce several consistent views of the edited objects or scenes.
In more detail, consider an initial reconstruction of a 3D scene in terms of a mixture of Gaussians $\mathcal{G}$ (Sec. 3) and consider a camera trajectory $(\pi_{t})_{t=1}^{T}$ consisting of $T$ viewpoints. We first render the corresponding views $\mathcal{I}=(I_{t})_{t=1}^{t}$ where $I_{t}=\operatorname{Rend}(\mathcal{G},\pi_{t})$ is obtained by utilizing Eq. 1. Let $y$ be the textual prompt that specifies the desired edits and let $\mathcal{I}^{\prime}=(I^{\prime}_{t})_{t=1}^{T}$ be the images obtained by editing images $\mathcal{I}$ according to $y$. We can then update the 3D model $\mathcal{G}$ by fitting it to $\mathcal{I}^{\prime}$ according to
@@
The key question is how to obtain the edited views $\mathcal{I}^{\prime}$. A standard approach is to apply a diffusion-based image editor $\mathcal{E}$, such as IP2P [107] to the individual views, obtaining edited views $\mathcal{I}^{\prime}$. However, each of these views would then be drawn independently from the distribution $I^{\prime}_{t}\sim p(I^{\prime}_{t}\mid I_{t},y).$ However, even though the initial views $I_{t}$ are consistent by construction, the edited views would not be as this method disregards their statistical dependency. As a result, updating the 3D model $\mathcal{G}$ according to Eq. 4 would fuse inconsistent views, yielding a blurry outcome. We thus seek to modify the image editor to account for the mutual dependency of the views and approximately draw samples from the joint distribution $p(I^{\prime}_{1},\dots,I^{\prime}_{t}\mid I_{1},\dots,I_{T},y)$.
In summary, we divide the 3D editing process into two parts: (a) multi-view consistent editing with epipolar constraints, and (b) 3D reconstruction from edited images, which we detail next.

### v2

@@ Abstract
We consider the problem of editing 3D objects and scenes based on open-ended language instructions. A common approach to this problem is to use a 2D image generator or editor to guide the 3D editing process, obviating the need for 3D data. However, this process is often inefficient due to the need for iterative updates of costly 3D representations, such as neural radiance fields, either through individual view edits or score distillation sampling. A major disadvantage of this approach is the slow convergence caused by aggregating inconsistent information across views, as the guidance from 2D models is not multi-view consistent. We thus introduce the Direct Gaussian Editor (DGE), a method that addresses these issues in two stages. First, we modify a given high-quality image editor like InstructPix2Pix to be multi-view consistent. To do so, we propose a training-free approach that integrates cues from the 3D geometry of the underlying scene. Second, given a multi-view consistent edited sequence of images, we directly and efficiently optimize the 3D representation, which is based on 3D Gaussian Splatting. Because it avoids incremental and iterative edits, DGE is significantly more accurate and efficient than existing approaches and offers additional benefits, such as enabling selective editing of parts of the scene.
@@ 1 Introduction
Recent breakthroughs in 2D and 3D generation [73, 28, 70, 56] have significantly advanced AI-based content creation and editing. This progress is particularly impactful for 3D content, which traditionally requires highly specialized skills and years of experience. Consequently, these innovations are unlocking new creative possibilities for artists and non-professional users alike.
@@
In this work, we consider the problem of editing 3D models based on textual instructions. Recent progress in this area is largely due to the introduction of new radiance field representations for objects and scenes, such as NeRFs [61], which are both high-quality and flexible. Because they are differentiable, radiance fields integrate well with applications of machine learning to image generation. A notable example is InstructPix2Pix (IP2P) [4], a denoising diffusion model [78, 73] that can edit an image based on textual instructions [4]. A significant benefit of differentiable 3D representations, when combined with image generation models, is that they eliminate the need for 3D annotated data. As a result, several works have repurposed such models to edit 3D content instead, primarily in two ways. The first was pioneered by InstructNeRF2NeRF [25], which utilizes IP2P to alternate between editing a rendered view of the 3D model and updating the latter by training on the edited images. The second approach [112, 34, 75] is to update the model in a distillation-like fashion [70]. Both approaches rely on iterative mechanisms to incorporate the edits into the 3D model, so that a single edit often requires several minutes, or even hours, to complete.
@@
In this paper, we introduce the Direct Gaussian Editor (DGE), a 3D editor that addresses the shortcomings of previous attempts. We design our method with three goals in mind: (i) high fidelity, (ii) high efficiency, and (iii) selective editing (i.e., editing only a specific part of the scene). To achieve these goals, we propose to change both the 3D representation and the update mechanism.
@@
As a representation, we employ 3D Gaussian Splatting (GS) [37], a radiance field model that is notably orders of magnitude faster than NeRF and NeRF-adjacent models [9, 80, 7] for both rendering and gradient computation. Additionally, GS offers another important advantage: as an explicit 3D representation made of local 3D primitives, the Gaussians, it supports local edits easily and efficiently, as long as the relevant Gaussians can be identified. In practice, one can easily identify the Gaussians by fusing the output of a 2D segmenter [39] from several rendered views of the scene.
@@
While GS can significantly speed up the 3D editing process, it does not remove the bottleneck caused by the need to iterate several times between rendering, reconstruction, and evaluation of the underlying image-based diffusion model. The iterative methods used in prior works are slow because of the lack of view-consistent edits that could be used to update the 3D model coherently. Since 2D editors like IP2P are monocular and provide a distribution over all possible edits, the probability of drawing independently two or more consistent edits from different views is nearly zero. As a result, iterative dataset updates or distillation are ways of incrementally reaching a “multi-view” consensus with respect to the application of the monocular editor to such views.
@@
Our main contribution is a more efficient alternative to this slow iterative process. We propose a method to sample multi-view consistent edits such that the 3D model can be updated by directly fitting it to the edited views. Our approach is inspired by recent progress in video generation and editing and, in particular, methods that extend image generators to video without additional training [21, 94, 38, 6, 71]. The key insight is that multiple views of the 3D model can be interpreted as an orbital video of a static scene generated by a moving camera, making multi-view consistency analogous to temporal consistency in video editing. To achieve multi-view consistency we adopt the spatio-temporal attention mechanism used in the video editing literature and extend it with additional epipolar constraints, which we can enforce due to the 3D nature of our problem. This method ensures that edits are coherent across multiple views, improving the efficiency and effectiveness of the 3D editing process.
@@
Through qualitative and quantitative comparisons with prior works, we demonstrate two key advantages of DGE, even compared to recent works such as the GaussianEditor [13], which also uses Gaussian Splatting. First, as a direct editing approach, DGE results in a noticeable speed boost taking approximately 4 min for a single edit. Second, ensuring multi-view consistent editing in the image space significantly simplifies the process of consolidating edits from various views into the 3D model. This is reflected in both the number of updates required for convergence and the higher fidelity of the final result.
@@ 2 Related Work
Due to the lack of large-scale training data to learn 3D editors directly, most such editors build on top of existing 2D image editors instead, so we discuss those first. GLIDE [66] controls 2D image generation and editing using CLIP features [72]. Methods like [19, 42] and DreamBooth [74] consider the problem of personalizing an image; GLIGEN [50] and others [8, 11, 17] perform layout control, while ControlNet [106] considers additional forms of control such as scribbles or depth maps. DragDiffusion [76] and Drag-a-Video [82] edit images and videos by dragging. Others [57, 64, 2, 85, 68, 36, 27] cast image editing as image-to-image translation. InstructPix2Pix (IP2P) [4] fine-tunes Stable Diffusion [73] with image-and-text conditioning on a large synthetic dataset of language-driven edits. The works of [105, 108] further improve InstructPix2Pix via manual labeling, while [62] aims at improving the localization of the edits.
2.0.2 Ad-hoc 3D Editing.
@@
Several authors have explored various types of inputs and controls to edit 3D objects. EditNeRF [53] updates shape and color in a radiance field based on user-provided scribbles, 3Designer [47] and SINE [1] based on a single edited view, SKED [59] based on 2D sketches, Editable-NeRF [109] based on keypoints and CoNeRF [35] based on attributes. The work of [53] also considers sketch-based editing but for category-level 3D generators. NeRF-Editing [101], NeuMesh [98] and NeRFShop [33] modify a radiance field based on meshes, NeuralEditor [10] based on point clouds and [97] based on cages. N3F [84], DFF [40], SPIn-NeRF [63], NeRF-in [52] and the work of [93] consider segmenting and removing objects from a radiance field. Control-NeRF [43] allows both removing and moving objects. Component-NeRF [51] and the works of [103, 99] consider compositional editing. Palette-NeRF [41], RecolorNeRF [22], ICE-NeRF [44] address recoloring and ARF [104], DeSRF [96], StylizedNeRF [31], SNeRF [65] and the work of [15, 30] style transfer. NeRFEditor [81] integrates GAN-based stylization in NeRF. Seal-3D [92] edits 3D scenes by learning interactive tools like brushes, deformation, and recoloring and Seal4D-NeRF [32] extends it to dynamic NeRF models. SceNeRFlow [83] allows for dynamic edits using DensePose for correspondence estimation [24].
2.0.3 Language-driven 3D Editing.
@@
Closer to our work, several authors have considered open-vocabulary or text-guided 3D editing. Some authors use vision-language models based on CLIP [72, 46, 48] to edit or stylize a radiance field globally [88, 89, 58, 45] or locally [77, 40, 90, 23, 110]. TextDeformer [20] focuses on manipulating only the shape of objects. Some like AvatarCLIP [29] consider animations as well. Most recent works have shifted towards instruction-guided editing employing diffusion-based image generators like Stable Diffusion [73] or editors like Instruct-Pix2Pix [4]. There exist two main editing mechanisms among them. Instruct-NeRF2NeRF [25], InstructP2P [95], ProteusNeRF [87], Edit-DiffNeRF [100], and GaussianEditor [13, 91] repeatedly edit rendered views of the 3D object and update the 3D model accordingly, a process which is referred to as iterative dataset updates. Instead, Instruct3Dto3D [34], DreamEditor [112], Vox-E [75], ED-NeRF [67], FocalDreamer [49], Progressive3D [14] and [102, 111] update the 3D model using score distillation sampling (SDS). While these works rely on costly, scene-specific optimization of radiance fields, SHAP-EDITOR [12] learns a fast, feed-forward editor in the latent space of generative models but requires retraining for each new set of instructions. Our editor is optimization-based, but significantly reduces the number of iterations required, up to a single one.
Closely related to our work, ViCa-NeRF [16] focuses on multi-view consistent editing. However, different from our method, which takes inspiration from video models, ViCA-NeRF leverages the 3D model’s depth information to project features from key views into others using a blending module. As a result, it cannot handle edits that change the geometry of the original 3D shape.
@@
Furthermore, methods like Vox-E [75] and FocalDreamer [49] consider local 3D editing, but struggle to handle both local and global ones; in contrast, our DGE allows precise editing of small or large regions of the 3D scene.
@@ 3 Preliminaries
3.0.1 Gaussian Radiance Fields.
A radiance field is a pair of functions $\sigma:\mathbb{R}^{3}\rightarrow\mathbb{R}^{+}$ and $c:\mathbb{R}^{3}\times\mathbb{S}^{2}\rightarrow\mathbb{R}^{3}$ mapping 3D points $\boldsymbol{x}\in\mathbb{R}^{3}$ to opacities $\sigma(\boldsymbol{x})$ and directional colors $c(\boldsymbol{x},\boldsymbol{\nu})$, where $\boldsymbol{\nu}\in\mathbb{S}^{2}$ is a unit vector expressing the viewing direction. An image $I$ of the radiance field is obtained via the emission-absorption equation:
@@
Prior works have explored several representations for these functions, including MLPs [61], voxel grids [80] and low-rank factorizations of the latter [9, 7]. Gaussian Splatting (GS) [37] proposes a particularly efficient representation as a mixture of Gaussians $\mathcal{G}=\{(\sigma_{i},\boldsymbol{\mu}_{i},\Sigma_{i},c_{i})\}_{i=1}^{G}$, where $\sigma_{i}\geq 0$ is the opacity, $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ is the mean, $\Sigma_{i}\in\mathbb{R}^{3\times 3}$ is the covariance matrix, and $c_{i}:\mathbb{S}\rightarrow\mathbb{R}^{3}$ the directional color of each Gaussian. The Gaussian functions are given by:
@@
3.0.2 Diffusion-based Generators and Editors.
A diffusion-based image generator is a model that allows one to sample image $I\in\mathbb{R}^{3\times H\times W}$ from a conditional distribution $p(I\mid y)$. Here the condition $y$ is often a textual prompt, but it can also be an image or a combination of both. Let $I_{k}=\sqrt{1-\beta_{k}^{2}}I+\beta_{k}\epsilon$ be a noised version of the image $I$ one wishes to sample, where $\beta$ is a sequence monotonically increasing from 0 to 1 and $\epsilon\sim\mathcal{N}$ is normal i.i.d. noise. The diffusion model consists of a denoising neural network $\Phi$ that, given the noisy image $I_{k}$, the noise level $k$ and the conditioning $y$, estimates the noise $\hat{\epsilon}=\Phi(I,y,k).$ The image $I$ is sampled by starting from pure noise ($\beta_{T}=1$) followed by iterative denoising of the signal until $I=I_{0}$ is obtained [28, 78, 86].
Several authors [57, 27, 68] have suggested using the denoising network of an off-the-shelf image generator such as Stable Diffusion (SD) [73, 69] to modify an existing image according to a prompt $y$ instead of generating it from scratch, thus editing it. However, this process is somewhat suboptimal; instead, we make use of editors specifically trained to apply such modifications. An example is IP2P [4], which also uses diffusion but directly implements conditional a distribution $p(I^{\prime}\mid I,y)$ where $I^{\prime}$ is the edited image, $I$ is the original image, and $y$ is a prompt describing how $I^{\prime}$ should be obtained from $I$. IP2P is in itself derived from the SD model by finetuning it to training triplets $(I^{\prime},I,y)$ which are automatically generated using Prompt-to-Prompt [27], a slow but training-free editing method.
@@ 4 Method
In this section, we describe our method for text-guided 3D scene editing — the Direct Gaussian Editor (DGE). Our main objective is 3D editing that is (i) faithful to the text instructions (high fidelity), (ii) fast and efficient, and (iii) partial to specific scene elements when desired. Next, we introduce our approach and discuss how these criteria are met.
@@ 4.1 Direct Gaussian Editor
DGE is motivated by the following key observation: given several images rendered from an existing 3D model, if one can obtain multi-view consistent edits in image space, then the 3D model can be updated directly and efficiently by fitting it to the edited images. This approach is an alternative to slow iterative techniques like SDS [70, 75, 49] and iterative dataset updates (IDU) [25, 13, 18] that work around inconsistent image-level edits. Thus, the main challenge for enabling direct 3D editing is to improve image-based editors to produce several consistent edits across a number of images.
In more detail, consider an initial reconstruction of a 3D scene as a mixture of Gaussians $\mathcal{G}$ (Sec. 3) and consider a camera trajectory $(\pi_{t})_{t=1}^{T}$ consisting of $T$ viewpoints. We first render the corresponding views $\mathcal{I}=(I_{t})_{t=1}^{T}$ where $I_{t}=\operatorname{Rend}(\mathcal{G},\pi_{t})$ is obtained with Eq. 1. Let $y$ be the textual prompt describing the desired edits and let $\mathcal{I}^{\prime}=(I^{\prime}_{t})_{t=1}^{T}$ be the images obtained by editing $\mathcal{I}$ as according to $y$. We can then update the 3D model $\mathcal{G}$ by fitting it to $\mathcal{I}^{\prime}$ as
@@
The key question is how to obtain the edited views $\mathcal{I}^{\prime}$. A standard approach is to apply a diffusion-based image editor $\mathcal{E}$, such as IP2P [95] to individual views, obtaining edited views $\mathcal{I}^{\prime}$. However, each of these views would then be drawn independently from the distribution $I^{\prime}_{t}\sim p(I^{\prime}_{t}\mid I_{t},y).$ Hence, even though the initial views $I_{t}$ are consistent by construction, the edited views would not be as this method disregards their statistical dependency. As a result, updating the 3D model $\mathcal{G}$ according to Eq. 4 would fuse inconsistent views, yielding a blurry outcome. We thus seek to modify the image editor to account for the mutual dependency of the views and approximately draw samples from the joint distribution $p(I^{\prime}_{1},\dots,I^{\prime}_{T}\mid I_{1},\dots,I_{T},y)$.
In summary, as shown in Figure 1, we divide the 3D editing process into two parts, detailed next: (a) multi-view consistent editing with epipolar constraints, and (b) the 3D reconstruction from edited images.